When the price of a product increases by 5% and the quantity demanded of the product changes from 1 500 to 1 200, what is the price elasticity of demand of the product? (Ignore the negative sign and round off to 2 decimal places)

1
Expert's answer
2021-08-11T09:12:52-0400

Price elasticity of demand= percentage change in quantity demanded divided by the percentage change in price.

Here, percentage change in quantity demanded "=\frac {1200-1500}{1500}\times 100" "=-20%" %

Price elasticity of demand "=\frac{-20}{5}=-4.00""= 4.00" 

The percentage change in quantity demanded is to a larger extent as compared to percentage change in price.